# Positive-Unlabeled Constraint Learning (PUCL) for Inferring Nonlinear Continuous Constraints Functions from Expert Demonstrations

### v1

@@ Abstract
Planning for a wide range of real-world robotic tasks necessitates to know and write all constraints. However, instances exist where these constraints are either unknown or challenging to specify accurately. A possible solution is to infer the unknown constraints from expert demonstration. This paper presents a novel Positive-Unlabeled Constraint Learning (PUCL) algorithm to infer a continuous arbitrary constraint function from demonstration, without requiring prior knowledge of the true constraint parameterization or environmental model as existing works. Within our framework, we treat all data in demonstrations as positive (feasible) data, and learn a control policy to generate potentially infeasible trajectories, which serve as unlabeled data. In each iteration, we first update the policy and then a two-step positive-unlabeled learning procedure is applied, where it first identifies reliable infeasible data using a distance metric, and secondly learns a binary feasibility classifier (i.e., constraint function) from the feasible demonstrations and reliable infeasible data. The proposed framework is flexible to learn complex-shaped constraint boundary and will not mistakenly classify demonstrations as infeasible as previous methods. The effectiveness of the proposed method is verified in three robotic tasks, using a networked policy or a dynamical system policy. It successfully infers and transfers the continuous nonlinear constraints and outperforms other baseline methods in terms of constraint accuracy and policy safety.
@@
Constraints inference, learning from demonstrations, inverse reinforcement learning, transfer learning
@@ I Introduction
Planning for many robotics and automation tasks requires explicit knowledge of constraints, which define what states or trajectories are allowed or must be avoided [1, 2]. Sometimes these constraints are initially unknown or hard to specify mathematically, especially when they are nonlinear, possess unknown shape of boundary or are inherent to user’s preference and experience. For example, human users may determine an implicit minimum distance of the robot to obstacles based on the material of the obstacle (glass or metal), the shape and density of obstacles and their own personal preference. The robot should be able to infer such a constraint as well as the desired minimum distance to achieve the task goal and meet human’s requirements. This means an explicit constraint should be inferred somehow, e.g., from existing human demonstration sets.
(a) Demonstrations
(b) Learned constraint
(c) Transferred policy
Figure 1: The task requires the robot to reach a target state (blue dot) while avoid going close to or over the cups. These unknown constraints can be translated into an infeasible region that must be inferred from a set of feasible paths demonstrated by an expert (in green). (a) Demonstrations (shown in green lines). (b) Learned constraint region (in red) (c) New transferred policy (in orange), generated by transferring the learned constraint to a similar task of reaching another target state (blue dot) from different starting points.
@@
IRL-based methods draw inspiration from the concepts and approaches of IRL and adapt them to learn the constraint function. [4] first formulate the constraint learning problem in the maximum likelihood inference framework. They introduce a Boltzmann policy model, where the likelihood of any feasible trajectory is assumed to be proportional to the exponential return of the trajectory, while the likelihood of any infeasible trajectory is 0. Then a greedy algorithm is proposed to add the smallest number of infeasible states that maximize the likelihood of the demonstrations. Further, [5] and [6] extend the maximum entropy framework from a deterministic setting into a stochastic setting, with soft or probabilistic constraints. [7] proposes an algorithm that learns a neural network policy via deep RL and generates high return trajectories with the policy network. Then the infeasible states are again added based on the maximum likelihood principle. [8] proposes Maximum Entropy Constraint Learning (MECL) algorithm. It not only approximates the policy with a network but also learns a constraint network, which is optimized by making a gradient ascent on the maximum likelihood objective function. More recently, [9] proposes a variational approach to infer the distribution of constraint and capture the epistemic uncertainty. However, despite the success of aforementioned IRL-based methods, they suffer from the two problems: theoretically, their maximum likelihood loss function optimizes all states visited by the policy to be infeasible, and at the same time optimizes states visited by the demonstrations to be feasible. If the demonstrations and policy overlap a lot, this loss will lead to conflicting updating, slowing the training and prone to classifying safe demonstrated states as infeasible. In addition to that, in practice their methods are mostly applied to recover linear constraints of a single state element, e.g. $x\geq-3$ [8, 9, 10], and, as we show later in the paper, perform relatively poorly in learning more complex constraints.
@@
In parallel, another line of research based on MIP has been developed. [3] first studies how to infer infeasible states from a demonstration set in a grid world. They assume that all possible trajectories that could earn higher rewards than the demonstration must be constrained in some way, or the expert could have passed those trajectories. In practice, they use hit-and-run sampling to obtain such higher-reward trajectories and solve an integer program to recover the constrained states in a grid world. This MIP method is soon extended to continuous state space and parameterized continuous constraint functions, where the constraint is represented as unions and intersections of axis-aligned box [11]. Later, the same author proposes to leverage the closed-form environmental model and the KKT optimality condition to infer the parameters of a known constraint parameterization [11], or a Gaussian Process constraint model [12]. However, these MPI-based methods either require the full knowledge of a differentiable model of the environment and agent [11, 12], which is not always available in practice , or can only learn constrained region expressed with unions and intersections of box [13].
@@
In summary, despite these impressive advances, it still remains to be shown how to learn continuous nonlinear constraint functions from a small sample of demonstrations, without full knowledge of true constraint parameterization and environmental model. To overcome the aforementioned challenge, we propose a two-step Positive-Unlabeled Constraint Learning (PUCL) method, inspired by a machine learning subarea Positive-Unlabeled (PU) learning [14] and extended from our previous work [15]. Within our framework, we treat all data in demonstrations as positive (feasible) dataset, and learn a policy to generate many potentially unsafe trajectories which serve as unlabeled data with unknown feasibility. Then a two-step PU learning technique is applied, where we first identify reliable infeasible data and secondly learn a binary feasibility classifier (i.e., constraint function) from the positive data and reliable infeasible data. The proposed method is flexible to learn complex-shaped constraint boundary without the need to know the true constraint parameterization. It also avoids the drawback of IRL-based method that mistakenly optimize and classify demonstrations as infeasible, thus improving the performance and accuracy.
The idea and approach of positive-unlabeled constraint learning was first explored in our previous work [15]. This work replaced the postprocessing approach in [15] with a more powerful two-step approach and solved more complex constraint learning tasks.
@@ II Preliminaries and Problem Statements
II-A Preliminaries
For a finite-horizon Markov decision process (MDP) [16, 17], states and actions are denoted by $s\in\mathcal{S}$ and $a\in\mathcal{A}$. $\gamma$ denotes the discount factor, and the potentially unknown real-valued reward is denoted by $r(s,a)$. A trajectory $\tau=\{s_{1},a_{1},\ldots,s_{T},a_{T}\}$ contains a sequence of state-action pairs in one episode. For the trajectory $\tau$, the total discounted reward (return) is defined as $r(\tau)=\sum_{t=1}^{T}\gamma^{t}r(s_{t},a_{t})$. A policy, the mapping between states and actions, is denoted by $a=\pi(s)$.
In this study, we focus on learning the Markovian state-action constraints, i.e., avoiding visiting some forbidden states (and actions) throughout the entire trajectory. Mathematically, we denote the true constraint set by $\mathcal{C^{*}}=\{(s,a)\in\mathcal{S}\times\mathcal{A}|(s,a)\text{ is truly infeasible}\}$. Thus, the true constraint is written as $(s,a)\notin\mathcal{C^{*}}$. This definition implicitly assumes that the studied constraint is (1) time-independent: the infeasible set remains unchanged for every timestep of every trajectory; (2) deterministic: a state-action pair is either truly feasible or truly infeasible, though the prediction model could output a probabilistic prediction of feasibility. (3) defined only on the current state (and potentially action), independent of previous states and actions. In addition to the unknown constraints, some known constraints could also be present in the task, e.g., goal state constraints and system dynamic constraints. For simplicity of explanation, all demonstrations and trajectories are assumed to satisfy the known constraints (trajectories violating known constraints are discarded and not used for inferring unknown constraint).
@@ II-B Problem Formulation
As in similar works [8, 3], we require that the reward function $r(s,a)$ is known in advance. This assumption is reasonable in many robotics applications, where the reward might be shortest path, shortest time to reach the goal. We emphasize that the unknown constraints are not indicated in the known reward function, i.e., the agent does not receive any penalty for violating the unknown constraints. Subsequently, the real task is written as a constrained optimization problem:
@@
where $r(\tau)$ represents the reward function of trajectory $\tau$, (1b) represents the unknown constraint to be learned.
The demonstration set $\mathcal{D}=\{\tau_{i}\}^{M_{d}}_{i=1}$ comprises trajectories demonstrated by a demonstrator, who may not necessarily be an expert but still to some extent solves the task. More concretely, we make the following assumption of the demonstrations optimality and feasibility:
@@
Any demonstration $\tau_{i}^{d}$ from $\mathcal{D}=\{\tau_{i}^{d}\}^{M_{d}}_{i=1}$ is feasible and $\delta$-sub-optimal solution to the problem 1:
@@
    (feasible): $\tau_{i}^{d}$ satisfies (1b) and all known constraints.
@@
    ($\delta$-sub-optimal): $(1-\delta)r(\tau^{*})\leq r(\tau_{i})$, where $\tau^{*}$ is the feasible optimal solution to the problem 1. $\delta\in[0,1)$ is a coefficient of sub-optimality. In practice it is treated as a hyper-parameter and specified according to the user’s confidence of demonstration optimality.
We formally define the constraint learning task as:
@@
The task of learning constraint from demonstration is to recover the unknown true constraint set $C^{*}$ in Problem 1 from the demonstration set $\mathcal{D}$, with the known reward function $r(\tau)$.
Note that in the rest of the paper we only consider the state constraints and omit actions, but the framework can be easily extended to state-action constraints by augmenting the state with the action to form an augmented state.
To represent and learn an arbitrary constraint set in continuous space, we approximate it with a constraint network $\zeta_{\theta}(s)\in(0,1)$, which induces a constraint set $\mathcal{C}_{\theta}=\{s\in\mathcal{S}|\zeta_{\theta}(s)\leq d\}$, where $\theta$ is the network weights and $d=0.5$ is the threshold.
@@ III-A Motivation: Constraint Inference as Positive-Unlabeled Learning
As in many IRL-based and MBI-based works [8, 3], we infer the underlying constraint by contrasting the (safe) demonstration with some potentially unsafe trajectories generated by a certain policy. The training follows an iterative framework to incrementally learn the constraint. In each iteration, we first sample a set of high-reward trajectories $\mathcal{P}=\{\tau_{i}\}^{M_{p}}_{i=1}$ by performing the current policy $\pi_{\phi}$ (discussed later in III-D). Each sampled trajectory consists of sequential state-action pairs $\tau=\{s_{j},a_{j}\}_{j=1}^{N_{i}}$. Similar to [3], we speculate that the trajectory which wins higher reward than demonstration is potentially unsafe and cheat by visiting some forbidden states. However, it remains unclear which specific state(s) within the trajectory are unsafe. In other words, trajectory $\tau$ visits both truly feasible states and truly infeasible states but both remain unlabeled. In contrast, it is certain that all states on the demonstrated trajectories are labeled feasible by Assumption 1. Our goal is to classify each single state as feasible or infeasible by learning from a batch of fully labeled feasible samples and another batch of unlabeled mixed samples.
@@
PU learning methods for binary classification mainly divide into three categories: two-step methods, biased learning and class-prior-based methods [14]. The latter two methods both rely on an important assumption, Sampled Completely At Random (SCAR), which requires the demonstrations and policy to have a similar distribution in the truly feasible region. Our former paper [15] already explored the class-prior-based methods. However, the required SCAR assumption is usually unrealistic since the state distribution largely depends on the specific task and is hard to control. Violation of this assumption leads to performance degradation. In contrast, the two-step methods are based on a separability assumption, which requires the two classes can be strictly separated. Since we limit our discussion to deterministic constraint and a state is either truly feasible or infeasible, the separability assumption naturally holds. Therefore, we develop our methods based on the two-step framework [18, 19].
III-B Method: Positive-Unlabeled Constraint Learning
@@
Our two-step PU learning methods consists of the following steps: (1) identifying reliable infeasible states from unlabeled trajectory set $\mathcal{P}$, and (2) using supervised learning techniques to learn a classifier from the feasible demonstrations $\mathcal{D}$ and reliable infeasible state set (denoted as $\mathcal{R}$). In step 1, the identification of $\mathcal{R}$ starts from the intuition that the states in the unlabeled set $\mathcal{P}$ that are very different from all states in the demonstration set $\mathcal{D}$ are highly likely to be truly infeasible. We propose a kNN-like approach, which is straightforward to comprehend and implement, to identify the reliable infeasible set $\mathcal{R}$. The unlabeled states are first ranked according to their distance to the $k$ nearest feasible states from demonstration set. The unlabeled states whose distance is higher than a threshold are selected as reliable infeasible states. Mathematically, the ranking score for each unlabeled data $s\in\tau_{i}\in\mathcal{P}$ is calculated by the average distance:
$S_{\mathcal{D}}(s)=\frac{1}{k}\sum_{s^{\prime}\in kNN(s,\mathcal{D})}dis(s,s^{\prime})$ (2)
where $kNN(s,\mathcal{D})$ indicates the $k$-nearest neighbours of $s$ in the set $\mathcal{D}$, and $dis(s,s^{\prime})$ is a distance metric selected based on the specific task. In this work, this distance is always measured using Euclidean distance. Additionally, considering the range of each dimension of the state could be very different, each dimension is preferred to be normalized before computing distance and nearest neighbours. The reliable infeasible set $\mathcal{R}$ consists of all unlabeled data whose score exceeds a threshold $d_{r}$.
=$\mathcal{R}=\{s\in\mathcal{P}|S_{\mathcal{D}}(s)\geq d_{r}\}$ (3)
@@
In practice, since $\mathcal{D}$ contains consecutive state series from trajectories, a smaller $k$ around 1-5 is preferred. For $d_{r}$, it indicates user’s belief over the size of true infeasible area. A lower $d_{r}$ tends to identify a larger $\mathcal{R}$, leading to a more conservative constraint network. Conversely, a higher $d_{r}$ tends to find a smaller $\mathcal{R}$ and a more radical constraint function. In the extreme case of $d_{r}=0$, nearly any state not visited by the demonstrations is classified as infeasible. We admit selecting $d_{r}$ can be tricky, but such a parameter is generally inevitable since learning constraints is known to be an ill-posed problem with infinite valid solutions [8].
@@
However, in the case where the demonstration closely adheres to the true constraint boundary with a distance shorter than $d_{r}$, kNN-like metric may fail to learn an accurate boundary, as the reliable infeasible data are all at least $d_{r}$ far from the demonstrations. To adapt to this case, we additionally expand $\mathcal{R}$ by adding the state(s) from each trajectory in $\mathcal{P}$ that is closest to the original $\mathcal{R}$, since these states are likely to between the current boudnary and the true boudnary.
$\mathcal{R}\leftarrow\mathcal{R}\cup\{s|\exists\tau_{i}\in\mathcal{P},s=\arg\min_{s\in\tau_{i}}S_{\mathcal{R}}(s)\}$ (4)
@@
where $S_{\mathcal{R}}(s)$ is the score function sharing similar definition as $S_{\mathcal{D}}(s)$ in (2) but with a different set $\mathcal{P}$.
@@
In step 2, we directly train a neural network binary classifier $\zeta_{\theta}(s)$ from the reliable infeasible data $\mathcal{R}$, the memory buffer $\mathcal{M}$ (discussed later in III-C) and the (feasible) demonstrations data $\mathcal{D}$ with a standard cross entropy loss:
$\mathcal{L}(\theta)=-\frac{1}{N+M}[\sum^{N}_{s_{i}\sim\mathcal{D}}\log\zeta_{\theta}(s_{i})-\sum^{M}_{s_{j}\sim\mathcal{R}\cup\mathcal{M}}\log(1-\zeta_{\theta}(s_{j}))]$ (5)
After training, the learned constraint set is expressed as $\mathcal{C}_{\theta}=\{s\in\mathcal{S}|\zeta_{\theta}(s)\leq 0.5\}$.

### v2

# Positive-Unlabeled Constraint Learning for Inferring Nonlinear Continuous Constraints Functions from Expert Demonstrations

@@ Abstract
Planning for diverse real-world robotic tasks necessitates to know and write all constraints. However, instances exist where these constraints are either unknown or challenging to specify accurately. A possible solution is to infer the unknown constraints from expert demonstration. This paper presents a novel two-step Positive-Unlabeled Constraint Learning (PUCL) algorithm to infer a continuous constraint function from demonstrations, without requiring prior knowledge of the true constraint parameterization or environmental model as existing works. We treat all data in demonstrations as positive (feasible) data, and learn a control policy to generate potentially infeasible trajectories, which serve as unlabeled data. The proposed two-step learning framework first identifies reliable infeasible data using a distance metric, and secondly learns a binary feasibility classifier (i.e., constraint function) from the feasible demonstrations and reliable infeasible data. The proposed method is flexible to learn complex-shaped constraint boundary and will not mistakenly classify demonstrations as infeasible as previous methods. The effectiveness of the proposed method is verified in four constrained environments, using a networked policy or a dynamical system policy. It successfully infers the continuous nonlinear constraints and outperforms other baseline methods in terms of constraint accuracy and policy safety. This work has been published in IEEE Robotics and Automation Letters (RA-L). Please refer to the final version at DOI:10.1109/LRA.2024.3522756.
@@
Learning from demonstration, reinforcement learning, transfer learning
@@ I Introduction
Planning for many robotics and automation tasks requires explicit knowledge of constraints, which define which states or trajectories are allowed or must be avoided [1, 2]. Sometimes these constraints are initially unknown or hard to specify mathematically, especially when they are nonlinear, possess unknown shape of boundary or are inherent to user’s preference and experience. For example, human users may determine an implicit minimum distance of the robot to obstacles based on the material of the obstacle (glass or metal), the shape and density of obstacles and their own personal preference. The robot should be able to infer such a constraint to achieve the task goal and meet human’s requirements. This means an explicit constraint should be inferred somehow, e.g., from existing human demonstration sets.
Figure 1: Learning an obstacle-avoidance constraint from demonstrations with proposed method. The task requires the robot to reach a target state while avoiding going close to or over the cups. These unknown constraints can be translated into an infeasible region that must be inferred from demonstrations. The top row illustrates one obstacle configuration, while the bottom row shows another. In each row, the left image displays the expert demonstrations (in green), the middle image shows the learned constraint region (in red), and the right image illustrates the learned policy, which is trained from the learned constraint and tested on a set of shifted goal states and starting points.
@@
IRL-based methods draw inspiration from the concepts and approaches of IRL and adapt them to learn the constraint function. [4] first formulate the constraint learning problem in the maximum likelihood inference framework. They introduce a Boltzmann policy model, where the likelihood of any feasible trajectory is assumed to be proportional to the exponential return of the trajectory, while the likelihood of any infeasible trajectory is 0. Then a greedy algorithm is proposed to add the smallest number of infeasible states that maximize the likelihood of the demonstrations. This framework has the advantage of being able to work with sub-optimal demonstrations. Further, [5] and [6] extend the maximum likelihood framework from a deterministic setting into a stochastic setting, with soft or probabilistic constraints. [7] proposes to learn a neural network policy via deep RL and identify discrete infeasible states with the policy network. All aforementioned methods only extract constraint from a pre-defined finite constraint set. To learn an arbitrary continuous constraint function, [8] proposes Maximum Entropy Constraint Learning (MECL) algorithm. It not only approximates the policy with a network but also learns a constraint network, which is optimized by making a gradient ascent on the maximum likelihood objective function. More recently, [9] proposes a variational approach to infer the distribution of constraint and capture the epistemic uncertainty. However, despite that [8, 9] can learn a continuous constraint function, they have following shortcomings: theoretically, their loss function optimizes over all the states visited by the policy to be infeasible, and states visited by the demonstrations to be feasible. If demonstrations and policy overlap, this loss will lead to conflicts during the update step, slowing the training time and prone to classifying safe demonstrated states as infeasible. In addition to that, in practice [8, 9, 10] are used to recover linear constraints, and on a unidimensional state, e.g. learn a constraint of the type $x\geq-3$, and, as we show later in the paper, perform relatively poorly when learning nonlinear constraints.
@@
In parallel, another line of research based on MIP has been developed. [3] first studies how to infer infeasible states from a demonstration set in a grid world. They assume that all possible trajectories that could earn higher rewards than the demonstration must be constrained in some way, or the expert could have passed those trajectories. In practice, they use hit-and-run sampling to obtain such higher-reward trajectories and solve an integer program to recover the constrained states in a grid world. This MIP method is soon extended to continuous state space and parameterized continuous constraint functions, where the constraint is represented as unions and intersections of axis-aligned boxes [11]. Later, the same authors propose to leverage the closed-form environmental model and the KKT optimality condition to infer the parameters of a known constraint parameterization [11], or a Gaussian Process constraint model [12]. However, these MIP-based methods either require the full knowledge of a differentiable model of the environment and agent [11, 12], which is not always available in practice , or can only learn constrained region expressed with unions and intersections of box [13].
@@
In summary, despite these impressive advances, it still remains to be shown how to learn continuous nonlinear constraint functions, without full knowledge of true constraint parameterization and environmental model. To overcome the aforementioned challenge, we propose a two-step Positive-Unlabeled Constraint Learning (PUCL) method, inspired by a machine learning subarea Positive-Unlabeled (PU) learning [14]. Within our framework, we treat all data in demonstrations as positive (feasible) dataset, and learn a policy to generate many potentially unsafe trajectories which serve as unlabeled data with unknown feasibility. Then a two-step PU learning technique is applied, where we first identify reliable infeasible data and secondly learn a binary feasibility classifier (i.e., constraint function) from the positive data and reliable infeasible data. The proposed method is flexible to learn complex-shaped constraint boundary without the need to know the true constraint parameterization. It also avoids the drawback of some IRL-based methods that mistakenly optimize and classify demonstrations as infeasible, thus improving the performance and accuracy.
@@ II Preliminaries and Problem Statements
II-A Preliminaries and Notations
We consider a finite-horizon deterministic Markov decision process (MDP) [15], where the states and actions are denoted by $s\in\mathcal{S}$ and $a\in\mathcal{A}$. $\gamma$ denotes the discount factor, and the real-valued reward is denoted by $r(s,a)$. A trajectory $\tau=\{s_{1},a_{1},\ldots,s_{T},a_{T}\}$ contains a sequence of state-action pairs in one episode. For the trajectory $\tau$, the total discounted reward (return) is defined as $r(\tau)=\sum_{t=1}^{T}\gamma^{t}r(s_{t},a_{t})$. A policy, the mapping between states and actions, is denoted by $a=\pi(s)$.
In this study, we focus on learning the Markovian state-action constraints, i.e., avoiding visiting some forbidden states (and/or actions) throughout the entire trajectory. Mathematically, we denote the true constraint set by $\mathcal{C^{*}}=\{(s,a)\in\mathcal{S}\times\mathcal{A}|(s,a)\text{ is truly infeasible}\}$. Thus, the true constraint is written as $(s,a)\notin\mathcal{C^{*}}$. This definition implicitly assumes that the studied constraint is (1) time-independent; (2) deterministic: a state-action pair is either truly feasible or infeasible. In addition to the unknown constraints, some known constraints could also be present in the task, e.g., goal state constraints and system dynamic constraints. For simplicity of explanation, all demonstrations and trajectories are assumed to satisfy the known constraints (trajectories violating known constraints are discarded and not used for inferring the unknown constraint).
We will approximate the true constraint set with a constraint network $\zeta_{\theta}$, with $\theta$ as the parameter. To simplify the notation and allow for more flexibility, we first introduce a generalized state $\bar{s}\equiv\psi(s,a)$ as the input to the network, where $\psi(s,a)$ is a user-defined feature function that can be customized to incorporate prior knowledge about the constraint structure. For example, if the constraint is known to depend solely on the state (or actions), one could define $\psi(s,a)=s$ (or $a$). Readers should keep in mind that, in the following sections, we use the generalized state $\bar{s}$ as shorthand for $\psi(s,a)$. Following this notation, the constraint network is written as $\zeta_{\theta}(\bar{s})\in[0,1]$, which induces a constraint set $\mathcal{C}_{\theta}=\{(s,a)|\zeta_{\theta}(\bar{s})\leq d\}$ with $d=0.5$ being the decision threshold.
The demonstration set $\mathcal{D}_{\tau}=\{\tau_{i}\}^{M_{d}}_{i=1}$ comprises trajectories demonstrated by a demonstrator. For convenience, we also introduce the set $\mathcal{D}=\{\bar{s}|(s,a)\in\tau_{i}\in\mathcal{D}_{\tau}\}$ consisting of generalized states from all the demonstrated trajectories. Similarly, we define the unlabeled data set $\mathcal{P}_{\tau}=\{\tau_{i}\}^{M_{p}}_{i=1}$ and $\mathcal{P}=\{\bar{s}|(s,a)\in\tau_{i}\in\mathcal{P}_{\tau}\}$.
@@ II-B Problem Formulation
As in similar works [8, 3], we require that the reward function to be known in advance. This assumption is reasonable in many robotics applications, where the reward might be shortest path, shortest time to reach the goal. We emphasize that the unknown constraints are not indicated in the known reward function, i.e., there is no negative reward for violating the unknown constraints. The real task is written as a constrained optimization problem:
@@
where $r(\tau)$ represents the reward function of trajectory $\tau$. We make the following assumption of the demonstrations optimality and feasibility:
@@
Any demonstration $\tau_{i}\in\mathcal{D}_{\tau}$ is feasible and $\delta$-sub-optimal solution to the problem 1:
@@
    (feasible): $\tau_{i}$ satisfies (1b) and all known constraints.
@@
    ($\delta$-sub-optimal): $(1-\delta)r(\tau^{*})\leq r(\tau_{i})$, where $\tau^{*}$ is the feasible optimal solution to the problem 1. $\delta\in[0,1)$ is a coefficient of sub-optimality.
In practice $\delta$ is treated as a hyper-parameter and specified according to the user’s confidence of demonstration optimality. We formally define the constraint learning task as:
@@
The task of learning constraint from demonstration is to recover the unknown true constraint set $C^{*}$ in Problem 1 from the demonstration set $\mathcal{D}_{\tau}$, given the known reward function $r(\tau)$.
@@ III-A Motivation: Constraint Inference as Positive-Unlabeled Learning
As in many IRL-based and MIP-based works [8, 3], we infer the underlying constraint by contrasting the (safe) demonstration with some potentially unsafe trajectories generated by a certain policy. The training follows an iterative framework to incrementally learn the constraint. In each iteration, we first sample a set of high-reward trajectories $\mathcal{P}_{\tau}$ by performing the current policy $\pi_{\phi}$ (discussed later in III-D). Similar to [3], we speculate that the trajectory which wins a higher reward than demonstration is potentially unsafe and cheats by visiting some forbidden states. However, it remains unclear which specific state(s) within the trajectory are unsafe. In other words, trajectory visits both truly feasible states and truly infeasible states but all states remain unlabeled. In contrast, it is certain that all states on the demonstrated trajectories are labeled feasible by Assumption 1. Our goal is to classify each single state as feasible or infeasible by learning from a batch of fully labeled feasible states and another batch of unlabeled mixed states.
@@
PU learning methods for binary classification mainly divide into three categories: two-step methods, biased learning and weight-based methods [14]. The latter two methods both rely on an important assumption, Sampled Completely At Random (SCAR), which requires the demonstrations and policy to have a similar distribution in the truly feasible region. This SCAR assumption could be unrealistic and limits its application to only specific environments. In contrast, the two-step methods do not require SCAR assumption and only require the two classes can be strictly separated. Since we limit our discussion to deterministic constraint, i.e., a state is either truly feasible or infeasible, this separability assumption naturally holds. Therefore, this paper will develop a constraint learning method from a two-step framework [16, 17].
III-B Method
Our two-step PU learning method consists of the following steps: (1) identifying reliable infeasible data (denoted as $\mathcal{R}$) from unlabeled set $\mathcal{P}$, and (2) using supervised learning techniques to learn a classifier from the feasible demonstrations $\mathcal{D}$ and reliable infeasible dataset $\mathcal{R}$.
@@ III-B Method
In step 1, the identification of $\mathcal{R}$ starts from the intuition that data in the unlabeled set $\mathcal{P}$ that are very different from all data in the demonstration set $\mathcal{D}$ are highly likely to be truly infeasible. We propose a kNN-like distance-based approach, which is straightforward to comprehend and implement, to identify the reliable infeasible set $\mathcal{R}$. The unlabeled data are first ranked according to their distance to the $k$ nearest feasible data from demonstration set  [^1] Since $\mathcal{D}$ contains consecutive data points from trajectories, it is recommended to use a small $k$ around 1-3.. Mathematically, the ranking score for each unlabeled data $\bar{s}\in\mathcal{P}$ is calculated by the average distance:
$S_{\mathcal{D}}(\bar{s})=\frac{1}{k}\sum_{\bar{s}^{\prime}\in kNN(\bar{s},\mathcal{D})}dis(\bar{s},\bar{s}^{\prime})$ (2)
where $kNN(\bar{s},\mathcal{D})$ indicates the $k$-nearest neighbors of $\bar{s}$ in the set $\mathcal{D}$, and $dis(\bar{s},\bar{s}^{\prime})$ is a distance metric selected based on the specific task. Then, the unlabeled data whose scores are higher than a threshold $d_{r}$ are selected as reliable infeasible data:
=$\mathcal{R}=\{\bar{s}\in\mathcal{P}|S_{\mathcal{D}}(\bar{s})\geq d_{r}\}$ (3)
@@
Intuitively $d_{r}$ indicates user’s belief over the size of true infeasible region. A lower $d_{r}$ tends to identify a larger $\mathcal{R}$, leading to a more conservative constraint network. Conversely, a higher $d_{r}$ tends to find a smaller $\mathcal{R}$ and a more radical constraint function. In the extreme case of $d_{r}=0$, nearly any state not visited by the demonstrations is classified as infeasible. We admit selecting $d_{r}$ can be tricky, but such a parameter is generally inevitable since learning constraints is known to be an ill-posed problem with infinite valid solutions [8].
In the case where all dimensions of $\bar{s}$ are homogeneous, i.e., having similar physical meanings and units, $d_{r}$ will share these units and can be selected heuristically. For example, the first case in our experimental section learns an obstacle-avoidance constraint in a 3D position space and $d_{r}$ roughly indicates how far the demonstration is from the true obstacle.
While in some high-dimensional case where $\bar{s}$ contains heterogeneous dimensions, identifying $\mathcal{R}$ with distance-based metric and determining the corresponding $d_{r}$ are more tricky. We propose to merge $\mathcal{D}$ and $\mathcal{P}$, standardize each dimension of the combined dataset to have zero mean and unit variance, and then use the standardized data to compute the scores $S_{\mathcal{D}}(\bar{s})$. Based on these scores, we select a percentile X and identify the top X percent of the highest-scoring data in $\mathcal{P}$ as the reliable infeasible data.
@@
In practice, we also notice that when the demonstration closely adheres to the true constraint boundary at a short distance, kNN-like metric may fail to learn an accurate boundary, as the reliable infeasible data are all at least $d_{r}$ far from the demonstrations. To adapt to this case, we expand $\mathcal{R}$ by adding the state from each trajectory in $\mathcal{P}$ that is closest to the original $\mathcal{R}$, since these data are likely to lie between the current boundary and the true boundary.
$\mathcal{R}\leftarrow\mathcal{R}\cup(\cup_{\tau_{i}\in\mathcal{P}_{\tau}}{\arg\min_{(s,a)\in\tau_{i}}S_{\mathcal{R}}(\bar{s})})$ (4)
@@
where $S_{\mathcal{R}}(\bar{s})$ is the score function sharing similar definition as $S_{\mathcal{D}}(\bar{s})$ in (2) but with a different set $\mathcal{R}$. Note that this expansion is performed only once per iteration.
@@
In step 2, we train a neural network binary classifier $\zeta_{\theta}(\bar{s})$ from the reliable infeasible data $\mathcal{R}$, the memory buffer $\mathcal{M}$ (discussed later in III-C) and the (feasible) demonstrations data $\mathcal{D}$ with a standard cross entropy loss:
$\mathcal{L}(\theta)=-\frac{1}{N+M}[\sum^{N}_{\bar{s}_{i}\sim\mathcal{D}}\log\zeta_{\theta}(\bar{s}_{i})-\sum^{M}_{\bar{s}_{j}\sim\mathcal{R}\cup\mathcal{M}}\log(1-\zeta_{\theta}(\bar{s}_{j}))]$ (5)